6. (10 pts) A high-school teacher in a low-income urban school used cash incentives to encourage

student learning in his AP statistics class. 15 of his 61 students scored a 5 on the AP statistics

exam. Worldwide, the proportion of students who scored a 5 in 2010 was 0.15. Is this evidence

that the proportion of students who would score a 5 on the AP statistics exam when taught by the

teacher using cash incentives is higher than the worldwide proportion of 0.15?

a. State hypotheses, find the P-value, and give your conclusions in the context of the

problem. Do you have any reservations about using the z test for proportions for this

data?

b. Does this study provide evidence that cash incentives cause an increase in the proportion

of 5s on the AP statistics exam? Explain your answer.

7. (10 pts) The editors of a magazine have noticed that people seem to believe that a successful life

depends on having good friends. They would like to have a story about this and use a headline

such as “Most Adults Believe Friends Are Important for Success.” So they have commissioned a

survey to ask a random sample of adults whether a successful life depends on having good

friends. In a random sample of 1027 adults, 53% said yes. Should the editors go ahead and use

their headline? Why or why not. Use a statistical method for your answer.

8. (10 pts) American roulette wheels have 38 numbers, of which 18 are red. You are checking a

roulette wheel to see if red occurs with the right relative frequency. You observe 500 sins, in

which the ball lands on red 194 times. Carry out the four steps on page 399 in a test of the null

hypothesis that the probability of getting red is 18/38. Use

α

= 0.10.
Answer #1

6) NULL HYPOTHESIS H0:

ALTERNATIVE HYPOTHESIS Ha:

alpha=0.1

p= 15/61= 0.246

Z=p-P/sqrt(P*Q/n)

Z= 0.246-0.15/sqrt(0.15*0.85/61)

Z= 0.096/0.0457

Z=2.10

The P-Value is .017864. The result is significant because p < .10.

Since sample size is greater than 30 therefore we will use z test.

NOTE: As per guidelines I have done the first please repost the second. Thank You